OpenStudy (anonymous):

how to find minimum value of f(x, y)= 3x^4+3y^4-xy? should i make first derivative and make the equation equals 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so take the derivative once, and solve for when f' is 0, then take the derivative again and plug in the points you found for when f' was 0, if they are positive then you have a minimum point

OpenStudy (zzr0ck3r):

you need to take partials

OpenStudy (zzr0ck3r):

then you have a system of equations set to 0

OpenStudy (zzr0ck3r):

3x^4+3y^4-xy so partial with respect to x 12x^3-y=0 partial with respect to y 12y^2-x=0

OpenStudy (zzr0ck3r):

that should be y^3

OpenStudy (anonymous):

after partial derivative what should i do?

OpenStudy (zzr0ck3r):

x=cuberoot(y/12) 12y^3-cuberoot(y/12)=0 12y^3=cuberoot(y/12) what is the only solution to this?

OpenStudy (zzr0ck3r):

I think I am doing this right, sorry its been a bit.

OpenStudy (zzr0ck3r):

do you know the answer?

OpenStudy (zarkon):

There will be 3 solutions to the system of equations...thought not all will correspond to a min

OpenStudy (zarkon):

*though

OpenStudy (zzr0ck3r):

what am I doing wrong? I thought we set the partials to 0, then solve?

OpenStudy (zarkon):

that is correct

OpenStudy (anonymous):

i dont know the answer..

OpenStudy (zarkon):

just as a reminder...the equation \(x=x^3\) has 3 solutions ... -1,0,1

OpenStudy (zarkon):

for the same reasons you will get 3 solutions to your system

OpenStudy (anonymous):

still confused..